# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ 6 Learning in Recurrent Models
Biologically-inspired Methods.   Despite its extensive and successful exploitation in training procedures for state-of-the-art models, the Backpropagation algorithm (BP) has been also criticized about its lack of biological plausibility [50]. The main issue is represented by the way in which BP performs credit assignment (see Section 2), with backpropagation-based synaptic weights adaptations that are unlikely to occur in real neuronal cells. Several other issues such as weight transport, non-locality, global feedback pathway, are summarized in recent works [163, 197, 210]. BPTT inherits all the plausibility-related issues from BP, with the addition that the need to store neural activities for error computations is further exacerbated by the requirement to store them over time—a very different procedure w.r.t. the local message-passing structure happening in the brain [210]. Among several biologically plausible synaptic modification rules, such as Hebbian Learning [148], Millidge et al. [197] recently proved that Predictive Coding [300] approximates BP in arbitrary computational graphs, including RNNs, by exploiting only local and plausible rules. Predictive Coding revolves around the perspective that the brain operates as a probabilistic generative model, consistently formulating predictions about its surroundings and adjusting internal hypotheses in accordance with sensory clues, in line with the Bayesian brain hypothesis [210]. In [197], the variational inference problem to be solved is that of inferring activation values of each node in the computational graph (i.e., activation values of neural units) given the start nodes (the input data) and end nodes (target level). The neural computational graph is enriched by error neurons/units that encode the difference between the activity at a given layer of the network and the one which is predicted/generated from the parent nodes. In the case of RNNs, both the output $y_{t}$ and the states $x_{t}$ are augmented with additional error units that compute the mismatch between actual values of the corresponding variable and the prediction obtained from the variational inference. The final goal is the optimization of a functional known as the variational free energy, aiming at improving the generative model accuracy and minimizing model complexity. Inspired from the same notions, Local Representation Alignment (LRA) [211] aims at optimizing a related variational free energy by minimizing the discrepancy between neural activations and ad-hoc local targets. LRA does not assume an underlying probabilistic generative model, and associates each layer with a target activity vector, so that synaptic weights are modified to guide the layer activity towards these targets. LRA ensures that realistic targets are chosen, considering the possible representation of a layer. Mismatch signals are computed as the first derivative of a distance function, and the target representation for each layer is determined based on accumulated mismatch signals from other layers. The synaptic plasticity rules involve adjustments to connections between layers and error message passing pathways. The effectiveness of training recurrent models with LRA has been explored in [209], where the authors propose the Parallel Temporal Neural Coding Network (P-TNCN), a layerwise parallelizable recurrent model which avoids unfoldings over the temporal dimension. The reader can refer to the very recent survey [210] for a comprehensive overview on biologically plausible learning rules.
@@ 7 Benchmarks
Devising appropriate benchmarks for evaluating the wide spectrum of architectures for processing long sequences is a fundamental step. Indeed, looking back at the whole history of sequence processing, up to nowadays, there is a clear lack of a well-established and consistent consensus in what benchmarks to use in comparing different algorithms. Ideally, benchmarks should should create conditions to investigate its ability to preserve long term dependencies, providing long sequences and tasks in which old information might play an important role. However, the meaning of long sequence has changes over the years. While in the early era of vanilla RNNs and LSTMs a sequence was considered to be “long” when constituted by a number of steps in the order of few tens or hundreds, current models (Recurrent Transformers, SSMs) are capable of processing sequences composed by thousands to millions of tokens [29].
In this section we describe the most common benchmarks that have been tackled since the dawn of RNNs up to the more recent Long Range Arena benchmark [282], and our analysis is aimed at providing a comprehensive view on benchmarking with long sequences. We mostly focus on the benchmarks and datasets exploited in the large set of papers described in this survey: Table LABEL:tab:benchs summarizes the benchmarks (first column) and the scientific papers in which such benchmarks have been exploited (subsequent columns, each column containing the papers described in a different section of this survey). We categorized the various benchmarks depending on their application field, distinguishing among synthetic benchmarks, computer vision, natural language, time series, audio/speech, up to the recent Long Range Arena [282], which includes multiple sub-benchmarks.
@@ 7.1 Synthetic Benchmarks
Copy Task.   The Copy task [105] is a largely known benchmarks to evaluate the network ability to recall information observed several time steps in the past, asking the network to reproduce the “initial” tokens of a long sequence. Data consist of sequences of $L$ + 20 digits. The first ten tokens $(a_{0},a_{1},...,a_{9})$ are randomly chosen from $\{1,...,8\}$ and represent the sequence to be “remembered”. The subsequent $L$ tokens are populated by noise (usually simply modeled by the integer 00), and the last ten tokens are all equal to the repeat marker (usually the integer $9$), that serves as an trigger to tell the model that it is time to reproduce the initial 10 tokens. In fact, the goal of the recurrent model is to output the starting ten tokens $(a_{0},...,a_{9})$ in the correct order during the last 10 time steps, whenever the repeat marker 9 is presented. The random guessing baseline has loss $\text{log}(8)\simeq 2.08$. Training and test examples are both generated this way. Intuitively, temporal credit assignment is critical in this task. Several relevant works tackled this task [105, 5, 191, 10, 86, 320, 217]. Gu et al. [85] observed that Linear Time Invariant SSMs (Section 4), encompassing linear recurrence and global convolutions, can efficiently tackle this task by focusing solely on tracking time rather than comprehensively understanding the data. For instance, this could involve constructing a convolution kernel of the right length, as also pointed out by Romero et al. [244]. For this reason, it was proposed to exploit the Selective Copying task (originally introduced by Jing et al. [126] as the Denoise task). This task disrupts the straightforward approach of Copy by introducing random spacing between relevant tokens (i.e., relevant tokens are interleaved by a varying number of noise-like tokens, see Figure 18), eliminating trivial solutions. Consequently, this adjustment necessitates content-aware reasoning for the network to effectively memorize and filter out relevant tokens while disregarding irrelevant ones. Lately, the Selective Copy task has been used to asses Large Language models capability of retraining long range dependencies [85, 57], with sequences of $L=4096$, and a vocabulary of $16$ tokens. Some other variants, such as the Capacity task, have been proposed to evaluated the case of sliding windows [293].
@@
Induction Heads.   In-context learning abilities of Large Language Models (LLMs) [^21] In-context learning is usually defined as the ability to learn directly from examples within the context (the prompt) of the input sequence at inference time, and use information from the input to generate the right answer for the output [206]. have been recently hypothesized to emerge from the so-called induction head mechanism, that was tested by means of the synthetic Induction Heads task, proposed in [206] (Figure 18). In this task, models are required to perform copy and associative recall operations. Specifically, the models are tasked with recalling the token immediately following a designated special token (referring to Figure 18: the black token is the special token, the token to be retrieved is the blue one). When the special token is presented once again, the model must be capable to subsequently retrieve and output the token following it within the context (the blue token in Figure 18). The successful execution of this task implies that the model possesses the ability to generalize to longer sequences than those it was trained on. During training, a vocabulary size of $t$ tokens is exploited ($t=16$ in [85, 57], various $t$ are tested in [227]), with sequences of length $L=256$, composed by tokens randomly sampled from the vocabulary. Additionally, the location of the special token within the sequence is also randomly determined. The model generalization and extrapolation abilities are assessed at test time on sequences with increasing lengths, ranging from $642^{6}=$ 64 to $5762^{20}=1,048,$ 576. Induction Heads is exploited in several recent works [85, 57].
@@
Phonebook Look-Up.   In the recently introduced PhoneBook task [124], the model is asked to process a synthetic phone-book (e.g., each line looks like “John Powell: 609-323 $7$”) and to return the phone number when given a name. Various sizes of the phone-books are tested. The main intuition behind this task is to prove that despite the efficiency at inference-time, SSMs are somewhat limited when compared to Transformers on tasks that require copying from the input context, as assessed by [124, 57].
@@ 7.2 Computer Vision Benchmarks
MNIST.   The MNIST dataset [151] has been widely exploited to asses performances of sequence processing models. There are two common alternative setups/tasks: sequence classification and generative autoregressive image modeling. In the former, each MNIST digit ($28\times 28$ pixels) is decomposed into a sequential pattern and classification is performed once the whole sequence has been processed. In the scientific literature, different modalities have been explored [^22] This also depends on the model architecture that is being tested. An RNN processes vectorial patterns (pixel-by-pixel or row-by-row), while Transformer-based models generally follow the Visual Transformer (ViT) [60] approach of processing a sequence of image patches., i.e., processing each MNIST digit following the order of scanlines, one pixel at a time, starting at the top-left corner of the image, and ending at the bottom-right corner (resulting in a sequence with $L=784$)—or alternatively, row-by-row, where each row represents the model input at a certain time $t$, resulting in $L=28$). We refer to both these dataset variations as Sequential MNIST (sMNIST) [150]. An harder variant is permuted sequential MNIST (psMNIST), where the same fixed random pixel permutation is applied to each digit of the dataset, resulting in an increased time-delay between interdependent pixels. The datasets sMNIST and psMNIST are exploited in [157, 249, 150, 90, 271, 130], amongst others. Differently, sequential image generation is performed by predicting images in pixel-by-pixel manner [134]. Image completions and unconditional samples are also usually considered. The former takes as input a portion of an image (that can be considered as a prompt, and it is usually approached by Transformers architectures [134]) and proceeds in generating the remainder of the image. The latter comprises the task of generating images with no condition in any context (such as a prompt text or another image). Ororbia et al. [209] proposed a video-based generative modeling assessment, exploiting BouncingMNIST, BouncingNotMNIST and BouncingFashionMNIST in sequence prediction, zero-shot adaption and online continual learning setting. In BouncingMNIST, each sequence is a $20$-frame video consisting of two randomly chosen digits from MNIST, moving and bouncing around in a $64\times 64$ frame. Digits are placed at random initial locations within the overall patch, are assigned a velocity and move and bounce off edges of the overall frame. When digits occupy the same location, they occlude each other and overlap. In BouncingNotMNIST, fonts/glyphs from the NotMNIST dataset [^23] https://yaroslavvb.blogspot.com/2011/09/notmnist-dataset.html are exploited, with the same characteristics (e.g., dimensions, sequence length, etc.) of BouncingNotMNIST. The same holds for BouncingFashionMNIST, with exemplars taken from the FashionMNIST dataset [308]. Sequence prediction is tackled both with an encoder/decoder approach (first $10$ frames are provided as input to the encoder, while the decoder has to generate the next $10$ frames) or also by training on input sequences of $10$ frames to reconstruct those $10$ frames as well as predict $10$ other future frames [275] (a recurrent architecture processes image features extracted by a convolutional model). In zero shot-adaptation, a model that is trained on one dataset (e.g., BouncingMNIST) is tested to generate samples from another set (e.g., BouncingNotMNIST). In the continual online learning scenario [177], as discussed in [209], models are trained on a stream composed of three of the aforementioned datasets, with a single pass over data (one epoch), mini-batches of size $1$ (i.e., pure online learning), which is a much more realistic setting when faced with infinite streams of patterns, with data coming from different tasks at different times. The models are one-shot trained on a concatenation of the three datasets. We frame these approaches under the genMNIST identifier in Table LABEL:tab:benchs.
@@
Action Recognition.   Video Activity Detection focuses on predicting actions at each time step within a video. Thus, unlike video classification, this task deals with long videos containing multiple overlapping activities, introducing a challenging problem due to the need of strongly exploiting and building an expressive temporal context. The recent TTM [251] (see Section 3) evaluated causally masked inference, providing activity predictions for incoming frames without considering future steps, in the Charades dataset [269], which comprises $9.8$k videos showcasing $157$ daily household activities, with $7.9$k training and $1.8$k validation clips. Videos can include multiple overlapping activities, requiring models to predict various activity classes per frame, considering interactions and longer temporal contexts. The average video length is $30$ seconds, creating a challenging scenario for temporal activity detection. The AVA v2.2 [91] dataset is also employed to test spatio-temporal activity detection, and is composed by bounding box annotations of $80$ atomic visual actions in $430$ $15$-minute movie clips. Li et al. [157] leverage the NTU RGB+D dataset [264], which is a skeleton-based action recognition task, composed by $56880$ sequences belonging to $60$ action classes and $20$ frames each.

### v2

@@ 6 Learning in Recurrent Models
Biologically-inspired Methods.   Despite its extensive and successful exploitation in training procedures for state-of-the-art models, the Backpropagation algorithm (BP) has been also criticized about its lack of biological plausibility [229]. The main issue is represented by the way in which BP performs credit assignment (see Section 2), with backpropagation-based synaptic weights adaptations that are unlikely to occur in real neuronal cells. Several other issues such as weight transport, non-locality, global feedback pathway, are summarized in recent works [68, 230, 231]. BPTT inherits all the plausibility-related issues from BP, with the addition that the need to store neural activities for error computations is further exacerbated by the requirement to store them over time—a very different procedure w.r.t. the local message-passing structure happening in the brain [231]. Among several biologically plausible synaptic modification rules, such as Hebbian Learning [232], Millidge et al. [230] recently proved that Predictive Coding [233] approximates BP in arbitrary computational graphs, including RNNs, by relying solely local and plausible rules. Predictive Coding revolves around the perspective that the brain operates as a probabilistic generative model, consistently formulating predictions about its surroundings and adjusting internal hypotheses in accordance with sensory data, in line with the Bayesian brain hypothesis [231]. In [230], the variational inference problem to be solved is that of inferring activation values of each node in the computational graph (i.e., activation values of neural units) given the start nodes (the input data) and target nodes. The neural computational graph is enriched by error neurons/units that encode the difference between the activity at a given layer of the network and that which is predicted or generated from its parent nodes. In the case of RNNs, both the output $y_{t}$ and the states $x_{t}$ are augmented with additional error units that compute the mismatch between the actual values of the corresponding variables and the prediction obtained via variational inference. The final goal is the optimization of a functional known as the variational free energy, aiming at improving the generative model accuracy and minimizing model complexity.
Inspired from the same notions, Local Representation Alignment (LRA) [234] aims at optimizing a related variational free energy by minimizing the discrepancy between neural activations and ad-hoc local targets. LRA does not assume an underlying probabilistic generative model, and associates each layer with a target activity vector, so that synaptic weights are modified to guide the layer activity towards these targets. LRA ensures that realistic targets are chosen, considering the possible representation of a layer. Mismatch signals are computed as the first derivative of a distance function, and the target representation for each layer is determined based on accumulated mismatch signals from other layers. The synaptic plasticity rules involve adjustments to connections between layers and error message passing pathways. The effectiveness of training recurrent models with LRA has been explored in [22], where the authors propose the Parallel Temporal Neural Coding Network (P-TNCN), a layerwise parallelizable recurrent model which avoids unfoldings over the temporal dimension. The reader can refer to the very recent survey [231] for a comprehensive overview on biologically plausible learning rules.
In summary, beyond architecture, recent work revisits training itself: modern approximations to RTRL, local credit-assignment schemes, lifted optimisation, and biologically inspired predictive-coding variants all aim to circumvent BPTT’s memory footprint and latency. These algorithmic developments complement the architectural trends surveyed earlier, completing the picture of next-generation recurrent sequence processing. Table 5 reports a comparative summary of the algorithms mentioned in this section, highlighting their key properties and complexities. In the next Section 7, we review the most popular benchmarks for evaluating the models surveyed so far.
Table 5: Comparative summary of two key features of training algorithms for recurrent models. We indicate if an algorithm is able to learn online on single components of each sequence, if the overall complexity (either spatial or temporal or both) is $\mathcal{O}(L)$ (assuming all other parameters fixed), and if it scales worse than $\mathcal{O}(d_{s}^{2})$ in term of its dependence from the hidden state size. See the paper text for more details.
Algorithm Online $\mathcal{O}(L)$? $\mathcal{O}(d_{s}^{k})$, $k>2$?
BPTT [15] No Yes No
RBP [196, 197, 198] No Yes No
RTRL [19] Yes No Yes
UORO [194] Yes No No
SnAp-1 [206] Yes No No
FPTT [58] Yes Yes No
Local Prop. [219] Yes No No
@@ 7 Benchmarks
Devising appropriate benchmarks for evaluating the wide spectrum of architectures for processing long sequences is a fundamental step. Indeed, looking back at the whole history of sequence processing, up to the present day, there is a clear lack of a well-established and consistent consensus on which benchmarks to use when comparing different algorithms. Ideally, benchmarks should create conditions to investigate a model’s ability to preserve long-term dependencies, providing long sequences and tasks in which old information might play an important role. However, the meaning of long sequence has changed over the years. While in the early era of vanilla RNNs and LSTMs, a sequence was considered to be “long” when it consisted a number of steps in the order of a few tens or hundreds, current models (Recurrent Transformers, SSMs) are capable of processing sequences composed of thousands to millions of tokens [124].
In this section , we describe the most common benchmarks that have been tackled since the dawn of RNNs up to the more recent Long Range Arena benchmark [35], and our analysis is aimed at providing a comprehensive view of benchmarking with long sequences. We mostly focus on the benchmarks and datasets used in the large set of papers described in this survey: Table LABEL:tab:benchs summarizes the benchmarks (first column) and the scientific papers in which such benchmarks have been used (subsequent columns, each column containing the papers described in a different section of this survey). We categorized the various benchmarks according to their application field, distinguishing among synthetic benchmarks, computer vision, natural language, time series, audio/speech, and the recent Long Range Arena [35], which includes multiple sub-benchmarks.
@@ 7.1 Synthetic Benchmarks
Copy Task.   The Copy task [28] is a largely known benchmarks to evaluate the network ability to recall information observed several time steps in the past, asking the network to reproduce the “initial” tokens of a long sequence. Data consist of sequences of $L$ + 20 digits. The first ten tokens $(a_{0},a_{1},...,a_{9})$ are randomly chosen from $\{1,...,8\}$ and represent the sequence to be “remembered”. The subsequent $L$ tokens are populated by noise (usually simply modeled by the integer 0), and the last ten tokens are all equal to the repeat marker (usually the integer $9$), that serves as an trigger to tell the model that it is time to reproduce the initial 10 tokens. In fact, the goal of the recurrent model is to output the starting ten tokens $(a_{0},...,a_{9})$ in the correct order during the last 10 time steps, whenever the repeat marker 9 is presented. The random guessing baseline has loss $\text{log}(8)\simeq 2.08$. Training and test examples are both generated this way. Intuitively, temporal credit assignment is critical in this task. Several relevant works tackled this task [28, 174, 206, 104, 39, 212, 223]. Gu et al. [146] observed that Linear Time Invariant SSMs (Section 4), encompassing linear recurrence and global convolutions, can efficiently tackle this task by focusing solely on tracking time rather than comprehensively understanding the data. For instance, this could involve constructing a convolution kernel of the right length, as also pointed out by Romero et al. [236]. For this reason, it was proposed to exploit the Selective Copying task (originally introduced by Jing et al. [237] as the Denoise task). This task disrupts the straightforward approach of Copy by introducing random spacing between relevant tokens (i.e., relevant tokens are interleaved by a varying number of noise-like tokens, see Figure 18), eliminating trivial solutions. Consequently, this adjustment necessitates content-aware reasoning for the network to effectively memorize and filter out relevant tokens while disregarding irrelevant ones. Lately, the Selective Copy task has been used to asses Large Language models capability of retraining long range dependencies [146, 44], with sequences of $L=4096$, and a vocabulary of $16$ tokens. Some other variants, such as the Capacity task, have been proposed to evaluated the case of sliding windows [30].
@@
Induction Heads.   In-context learning abilities of Large Language Models (LLMs) [^22] In-context learning is usually defined as the ability to learn directly from examples within the context (the prompt) of the input sequence at inference time, and use information from the input to generate the right answer for the output [238]. have been recently hypothesized to emerge from the so-called induction head mechanism, that was tested by means of the synthetic Induction Heads task, proposed in [238] (Figure 18). In this task, models are required to perform copy and associative recall operations. Specifically, the models are tasked with recalling the token immediately following a designated special token (referring to Figure 18: the black token is the special token, the token to be retrieved is the blue one). When the special token is presented once again, the model must be capable to subsequently retrieve and output the token following it within the context (the blue token in Figure 18). The successful execution of this task implies that the model possesses the ability to generalize to longer sequences than those it was trained on. During training, a vocabulary size of $t$ tokens is exploited ($t=16$ in [146, 44], various $t$ are tested in [74]), with sequences of length $L=256$, composed by tokens randomly sampled from the vocabulary. Additionally, the location of the special token within the sequence is also randomly determined. The model generalization and extrapolation abilities are assessed at test time on sequences with increasing lengths, ranging from $2^{6}=64$ to $2^{20}=1,048,576$. Induction Heads is exploited in several recent works [146, 44].
@@
Phonebook Look-Up.   In the recently introduced PhoneBook task [241], the model is asked to process a synthetic phone-book (e.g., each line looks like “John Powell: 609-323 7777”) and to return the phone number when given a name. Various sizes of the phone-books are tested. The main intuition behind this task is to prove that despite the efficiency at inference-time, SSMs are somewhat limited when compared to Transformers on tasks that require copying from the input context, as assessed by [241, 44].
@@ 7.2 Computer Vision Benchmarks
MNIST.   The MNIST dataset [242] has been widely exploited to asses performances of sequence processing models. There are two common alternative setups/tasks: sequence classification and generative autoregressive image modeling. In the former, each MNIST digit ($28\times 28$ pixels) is decomposed into a sequential pattern and classification is performed once the whole sequence has been processed. In the scientific literature, different modalities have been explored [^23] This also depends on the model architecture that is being tested. An RNN processes vectorial patterns (pixel-by-pixel or row-by-row), while Transformer-based models generally follow the Visual Transformer (ViT) [243] approach of processing a sequence of image patches., i.e., processing each MNIST digit following the order of scanlines, one pixel at a time, starting at the top-left corner of the image, and ending at the bottom-right corner (resulting in a sequence with L=784L=784)—or alternatively, row-by-row, where each row represents the model input at a certain time $t$, resulting in $L=28$). We refer to both these dataset variations as Sequential MNIST (sMNIST) [235]. An harder variant is permuted sequential MNIST (psMNIST), where the same fixed random pixel permutation is applied to each digit of the dataset, resulting in an increased time-delay between interdependent pixels. The datasets sMNIST and psMNIST are exploited in [31, 186, 235, 40, 42, 58], amongst others. Differently, sequential image generation is performed by predicting images in pixel-by-pixel manner [72]. Image completions and unconditional samples are also usually considered. The former takes as input a portion of an image (that can be considered as a prompt, and it is usually approached by Transformers architectures [72]) and proceeds in generating the remainder of the image. The latter comprises the task of generating images with no condition in any context (such as a prompt text or another image). Ororbia et al. [22] proposed a video-based generative modeling assessment, exploiting BouncingMNIST, BouncingNotMNIST and BouncingFashionMNIST in sequence prediction, zero-shot adaption and online continual learning setting. In BouncingMNIST, each sequence is a $20$-frame video consisting of two randomly chosen digits from MNIST, moving and bouncing around in a $64\times 64$ frame. Digits are placed at random initial locations within the overall patch, are assigned a velocity and move and bounce off edges of the overall frame. When digits occupy the same location, they occlude each other and overlap. In BouncingNotMNIST, fonts/glyphs from the NotMNIST dataset [^24] https://yaroslavvb.blogspot.com/2011/09/notmnist-dataset.html are exploited, with the same characteristics (e.g., dimensions, sequence length, etc.) of BouncingNotMNIST. The same holds for BouncingFashionMNIST, with exemplars taken from the FashionMNIST dataset [244]. Sequence prediction is tackled both with an encoder/decoder approach (first $10$ frames are provided as input to the encoder, while the decoder has to generate the next $10$ frames) and by training on input sequences of $10$ frames to reconstruct those $10$ frames as well as predict $10$ other future frames [245] (a recurrent architecture processes image features extracted by a convolutional model). In zero-shot adaptation, a model that is trained on one dataset (e.g., BouncingMNIST) is tested on its ability to generate samples from another set (e.g., BouncingNotMNIST). In the continual online learning scenario [46], as discussed in [22], models are trained on a stream composed of three of the aforementioned datasets, with a single pass over the , using mini-batches of size $1$ (i.e., pure online learning), which is a much more realistic setting when faced with infinite streams of patterns, where data arrives from different tasks at different times. The models are trained in a one-shot manner on a concatenation of the three datasets. We frame these approaches under the genMNIST identifier in Table LABEL:tab:benchs.
@@
Action Recognition.   Video Activity Detection focuses on predicting actions at each time step within a video. Thus, unlike video classification, this task deals with long videos containing multiple overlapping activities, introducing a challenging problem due to the need to strongly exploit and build an expressive temporal context. The recent TTM [115] (see Section 3) evaluated causally masked inference, providing activity predictions for incoming frames without considering future steps, in the Charades dataset [252], which contains 9.89.8k videos showcasing $157$ daily household activities, with 7.97.9k training and 1.81.8k validation clips. Videos can include multiple overlapping activities, requiring models to predict various activity classes per frame, taking into account interactions and longer temporal contexts. The average video length is $30$ seconds, creating a challenging scenario for temporal activity detection. The AVA v2.2 [253] dataset is also employed to test spatio-temporal activity detection, and is composed of bounding box annotations of $80$ atomic visual actions in $430$ $15$-minute movie clips. Li et al. [31] leverage the NTU RGB+D dataset [254], which is a skeleton-based action recognition task, comprising $56880$ sequences belonging to $60$ action classes and each containing $20$ frames.